# MultiBLiMP 1.0: A Massively Multilingual Benchmark of Linguistic Minimal Pairs

### v2

@@ 7.2 Impact of Model Properties
The fitted model yields an $R^{2}$ of 0.936. Model size has a significantly positive impact on performance ($\beta=4.30$; $t=15.88$). Post-training, however, has significantly negative impact on performance ($\beta=-3.29$; $t=-6.47$). This finding is in line with Chirkova and Nikoulina (2024), who report that instruction tuning (a stage of post-training) can have a negative effect on multilingual fluency in LLMs. [^8] We experimented with different chat templates for the chat version of the Gemma3 models, but did not observe any improvement. We leave more elaborate ‘prompt engineering’ open for future work. We visualize this trend in Figure 5 for the four largest model families: performance increases with size, and post-trained models consistently underperform their base version.
@@ 7.3 Cross-Model Comparisons
We consider the impact of language-specific post- and pre-training in Fig. 6a) and 6b), comparing the chat version of Llama3-8B to Aya-8B and the base version of Llama3-8B to EuroLLM-9B. We have highlighted the specific languages on which Aya has been post-trained, and EuroLLM has been pre-trained. Languages outside this selection do not appear to benefit from cross-lingual transfer, as both Aya and EuroLLM perform worse than Llama3-8B in the languages they were not trained on, and McNemar testing shows Llama3-8B is significantly better than Aya for 43 languages, and better than EuroLLM for 31 languages. This suggests that language-specific pre-training has a stronger effect on linguistic ability than post-training: EuroLLM significantly outperforms Llama3-8B on 19 of its 32 languages (59%), whereas Aya only outperforms it on 8 out of its 18 languages (44%). Further research in a setting with better control over the training data is needed to identify more precisely the data properties that drive multilingual grammar acquisition.
Next, we consider the impact of model size in Figure 6c), by comparing the 8B model to the 70B model. The average performance goes up from 86.9% to 90.2% for this model, driven by gradual improvements: no language improves by more than 10%,, but almost all languages improve to some degree. We also conduct a McNemar test between the two models, with $p<0.05$, resulting in Llama3-70B being significantly better for 48 languages, and none the other way around.
Finally, we look at the impact of monolingual vs multilingual language modelling in Figure 6d. Goldfish models significantly outperform Llama3-8B for 39 out of 70 languages, whereas Llama only does so for 7 languages. This warrants a more detailed investigation in future work, to assess if this difference is solely driven by differences in data frequency and data quality, or by the multilingual LLM having to handle all these languages in a shared representation space (Chang et al., 2024a).
@@ Results
As shown in Fig. 7, the agreement type coefficients show that person agreement yields high $\Delta$ scores (i.e. more confidence in detecting the grammatical sentence). This can be due to the fact that first and second person subjects tend to be closed-class pronouns, whereas gender and number agreement requires acquisition of open-class subject features. Surprisingly, subject-verb distance does not have a negative impact on $\Delta$: one might expect agreement to be more challenging for long-distance dependencies. An increase in sentence perplexity strongly decreases $\Delta$: the more surprised a model is by a sentence, the less likely it is to make the right grammaticality judgment.
@@
Another important factor with a negative effect is the subword delta: if the correct form of the verb is split into more subwords than the incorrect form, the model is more likely to make a wrong judgment. As we expected, incongruent attractors result in a decrease in $\Delta$, while congruent attractors boost it. Finally, language frequency has a positive effect on the performance of the LLMs, whereas for Goldfish (which controls for frequency) this is not the case. This is the only factor for which the Goldfish coefficient is opposite to the LLMs, for all other factors it is similar. A more detailed breakdown of impact of language frequency on Gemma3-27B is shown in Fig. 1, where we highlight the languages that over- and underperform with respect to a fitted logistic curve. An interesting question for future work would be to investigate the extent to which the variance in performance is driven by typological features.
@@ 8 Discussion and Conclusion
Based on the results of evaluating 17 LMs on MultiBLiMP 1.0 and analyzing the results, we formulate several recommendations for the training of future massively multilingual models. Based on the comparison between Aya and EuroLLM in fig. 6, we surmise that to boost specific focus languages, fine-tuning is not as effective as pre-training. We observed that the Goldfish models significantly outperform LLMs as large as 70B parameters on 14 out of 101  languages, many of them low-resource. This suggests that formal linguistic competence can greatly suffer when the training data of a particular language only forms a tiny part of the general mix—especially if that language is not closely related to the bulk of this mix. Thus, communities of languages underserved by current NLP technology may be best helped not by integrating their languages into LLM projects, but by more targeted regional LM initiatives. Our finding that the accuracy on MultiBLiMP 1.0 depended strongly on the difference in number of subwords between the correct and the incorrect inflection is in keeping with previous work (Rust et al., 2021, e.g.,) which suggested that performance differences between languages in multilingual LLMs are strongly driven by the space allocated to the language in the tokenizer. Taken together, these results call for new tokenization strategies to mitigate this issue (Mielke et al., 2021).
@@
The creation of MultiBLiMP 1.0 was only possible due to the existence of UD and UniMorph, both large resources created by a large number of annotators over many years. While these resources have become increasingly marginalised since the introduction of LLMs, we see this work as a prime example of their continued usefulness. The wealth of linguistic knowledge that is captured by these resources will continue to prove invaluable for informed linguistic evaluation of LLMs (Opitz et al., 2025).
@@
We see two major opportunities for future work. First, we plan to expand MultiBLiMP 1.0 to more constructions beyond agreement, investigating more diverse phenomena not attested in English and broadening the current language set. Second, this benchmark presents a unique opportunity for computational typology, due to the diverse set of languages included and the number of phenomena covered. We hope that MultiBLiMP will enable learnability studies across typologically diverse languages, bringing new insights into their linguistic structure, but also into the work that is needed to put them onto equal footing in language modelling.
@@ Acknowledgments
We would like to thank Amir Kargaran for providing the language frequencies in CommonCrawl. We further thank Ahmet Üstün, Charlotte Pouw, Jirui Qi, Omer Goldman, Kanishka Misra and Will Merrill, as well as the GroNLP group for helpful feedback. Jaap Jumelet and Arianna Bisazza were supported by NWO grant VI.Vidi.221C.009. Leonie Weissweiler was supported by a postdoctoral fellowship from the German Research Foundation (DFG). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-13403.
@@ B.1 Language Distribution
B.2 Significantly Best Models per Language
@@
Goldfish Albanian $99.2~{}(p=0.041)$
Buriat $91.3~{}(p=0.002)$
Catalan $97.7~{}(p=0.004)$
Erzya $73.7~{}(p=0.000)$
Faroese $99.6~{}(p=0.006)$
Galician $98.3~{}(p=0.001)$
Icelandic $98.4~{}(p=0.000)$
Ligurian $94.5~{}(p=0.000)$
Marathi $95.2~{}(p=0.001)$
Northern Kurdish $94.7~{}(p=0.033)$
Northern Sami $96.9~{}(p=0.000)$
Welsh $99.4~{}(p=0.000)$
Wolof $97.0~{}(p=0.000)$
Yakut $95.8~{}(p=0.029)$
Llama3-70B Church Slavonic $68.1~{}(p=0.009)$
Gothic $75.4~{}(p=0.000)$
Komi-Zyrian $77.2~{}(p=0.070)$
Latin $92.1~{}(p=0.026)$
Old French $81.8~{}(p=0.001)$
Old Russian $80.0~{}(p=0.024)$
Sanskrit $81.4~{}(p=0.000)$
aya-32b French $99.5~{}(p=0.063)$
gemma3-27b Bulgarian $99.2~{}(p=0.000)$
Gheg Albanian $80.4~{}(p=0.050)$
Polish $98.8~{}(p=0.002)$

### v3

@@ 7.2 Impact of Model Properties
The fitted model yields an $R^{2}$ of 0.936. Model size has a significantly positive impact on performance ($\beta=4.30$; $t=15.88$). Post-training, however, has significantly negative impact on performance ($\beta=-3.29$; t=6.47t=-6.47). This finding is in line with Chirkova and Nikoulina (2024), who report that instruction tuning (a stage of post-training) can have a negative effect on multilingual fluency in LLMs. [^8] We experimented with different chat templates for the chat version of the Gemma3 models, but did not observe any improvement. We leave more elaborate ‘prompt engineering’ open for future work. We illustrate this in Figure 5 for the four largest model families: performance increases with size, and post-trained models consistently underperform their base version.
@@ 7.3 Cross-Model Comparisons
We consider the impact of language-specific post- and pre-training in Fig.˜6a) and 6b), comparing the chat version of Llama3-8B to Aya-8B and the base version of Llama3-8B to EuroLLM-9B. We have highlighted the specific languages on which Aya has been post-trained, and EuroLLM has been pre-trained. Languages outside this selection do not appear to benefit from cross-lingual transfer, as both Aya and EuroLLM perform worse than Llama3-8B in the languages they were not trained on, and McNemar testing shows Llama3-8B is significantly better than Aya for 43 languages, and better than EuroLLM for 31 languages. This suggests that language-specific pre-training has a stronger effect on linguistic ability than post-training: EuroLLM significantly outperforms Llama3-8B on 19 of its 32 languages (59%), whereas Aya only outperforms it on 8 out of its 18 languages (44%). Further research in a setting with better control over the training data is needed to identify more precisely the data properties that drive multilingual grammar acquisition.
Next, we consider the impact of model size in Figure 6c), by comparing the 8B model to the 70B model. The average performance goes up from 86.9% to 90.2% for this model, driven by gradual improvements: no language improves by more than 10%, but almost all languages improve to some degree. We also conduct a McNemar test between the two models, with $p<0.05$, resulting in Llama3-70B being significantly better for 48 languages, and none the other way around.
Finally, we look at the impact of monolingual vs. multilingual language modelling in Figure 6d. Goldfish models significantly outperform Llama3-8B for 39 out of 70 languages, whereas Llama performs better only for 7 languages. This warrants a more detailed investigation in future work, to assess if this difference is solely driven by differences in data frequency and data quality, or by the multilingual LLM having to handle all these languages in a shared representation space (Chang et al., 2024a).
@@ Results
As shown in Fig.˜7, the agreement type coefficients show that person agreement yields high $\Delta$ scores (i.e. more confidence in detecting the grammatical sentence). This can be due to the fact that first and second person subjects tend to be closed-class pronouns, whereas gender and number agreement requires acquisition of open-class subject features. Surprisingly, subject-verb distance does not have a negative impact on $\Delta$: one might expect agreement to be more challenging for long-distance dependencies. An increase in sentence perplexity strongly decreases $\Delta$: the more surprised a model is by a sentence, the less likely it is to make the right grammaticality judgment.
@@
Another important factor with a negative effect is the subword delta: if the correct form of the verb is split into more subwords than the incorrect form, the model is more likely to make a wrong judgment. As we expected, incongruent attractors result in a decrease in $\Delta$, while congruent attractors boost it. Finally, language frequency has a positive effect on the performance of the LLMs, whereas for Goldfish (which controls for frequency) this is not the case. This is the only factor for which the Goldfish coefficient is opposite to the LLMs, for all other factors it is similar. A more detailed breakdown of impact of language frequency on Gemma3-27B is shown in Fig.˜1, where we highlight the languages that over- and underperform with respect to a fitted logistic curve. An interesting question for future work would be to investigate the extent to which the variance in performance is driven by typological features.
@@ 8 Discussion and Conclusion
Based on the results of evaluating 42 LMs on MultiBLiMP 1.0 and analysing the results, we formulate several recommendations for the training of future massively multilingual models. Based on the comparison between Aya and EuroLLM in fig.˜6, we surmise that to boost specific focus languages, fine-tuning is not as effective as pre-training. We observed that the Goldfish models significantly outperform LLMs as large as 70B parameters on 14 out of 101  languages, many of them low-resource. This suggests that formal linguistic competence can greatly suffer when the training data of a particular language only forms a tiny part of the general mix—especially if that language is not closely related to the bulk of this mix. Thus, communities of languages underserved by current NLP technology may be best helped not by integrating their languages into LLM projects, but by more targeted regional LM initiatives. Our finding that the accuracy on MultiBLiMP 1.0 depended strongly on the difference in number of subwords between the correct and the incorrect inflection is in keeping with previous work (Rust et al., 2021, i.a.,) which suggested that performance differences between languages in multilingual LLMs are strongly driven by the space allocated to the language in the tokenizer. Taken together, these results call for new tokenization strategies to mitigate this issue (Mielke et al., 2021).
@@
The creation of MultiBLiMP 1.0 was only possible due to the existence of UD and UniMorph, both large resources created by a large number of annotators over many years. While these resources have become increasingly marginalised since the introduction of LLMs, we see this work as a prime example of their continued usefulness. The wealth of linguistic knowledge that is captured by these resources will continue to prove invaluable for informed linguistic evaluation of LLMs (Opitz et al., 2024).
@@
We see two major opportunities for future work. First, we plan to expand MultiBLiMP 1.0 to more constructions beyond subject-verb agreement, investigating more diverse phenomena not attested in English and broadening the current language set. Second, this benchmark presents a unique opportunity for computational typology, due to the diverse set of languages included and the number of phenomena covered. We hope that MultiBLiMP will enable learnability studies across typologically diverse languages, bringing new insights into their linguistic structure, but also into the work that is needed to put them onto equal footing in language modelling.
@@ Acknowledgments
We would like to thank Amir Kargaran for providing the language frequencies in CommonCrawl. We further thank Ahmet Üstün, Charlotte Pouw, Jirui Qi, Omer Goldman, Kanishka Misra and Will Merrill, as well as the GroNLP and Edinburgh ILCC groups for helpful feedback. We thank the TACL reviewers for their valuable feedback. Jaap Jumelet and Arianna Bisazza were supported by NWO grant VI.Vidi.221C.009. Leonie Weissweiler was supported by a postdoctoral fellowship from the German Research Foundation (DFG, WE 7627/1-1). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-13403.
@@ B.1 Language Distribution
B.2 Significantly Best Model per Language
@@
Goldfish Albanian 99.2(p=0.041)99.2~(p=0.041)
Buriat 91.3(p=0.002)91.3~(p=0.002)
Catalan 97.7(p=0.004)97.7~(p=0.004)
Erzya 73.7(p=0.000)73.7~(p=0.000)
Faroese 99.6(p=0.006)99.6~(p=0.006)
Galician 98.3(p=0.001)98.3~(p=0.001)
Icelandic 98.4(p=0.000)98.4~(p=0.000)
Ligurian 94.5(p=0.000)94.5~(p=0.000)
Marathi 95.2(p=0.001)95.2~(p=0.001)
Northern Kurdish 94.7(p=0.033)94.7~(p=0.033)
Northern Sami 96.9(p=0.000)96.9~(p=0.000)
Welsh 99.4(p=0.000)99.4~(p=0.000)
Wolof 97.0(p=0.000)97.0~(p=0.000)
Yakut 95.8(p=0.029)95.8~(p=0.029)
[rgb]0.753,0.753,0.753[rgb]0.0, 0.0, 0.0 Llama3-70B Church Slavonic 68.1(p=0.009)68.1~(p=0.009)
Gothic 75.4(p=0.000)75.4~(p=0.000)
Komi-Zyrian 77.2(p=0.070)77.2~(p=0.070)
Latin 92.1(p=0.026)92.1~(p=0.026)
Old French 81.8(p=0.001)81.8~(p=0.001)
Old Russian 80.0(p=0.024)80.0~(p=0.024)
Sanskrit 81.4(p=0.000)81.4~(p=0.000)
[rgb]0.753,0.753,0.753[rgb]0.0, 0.0, 0.0 aya-32b French 99.5(p=0.063)99.5~(p=0.063)
[rgb]0.753,0.753,0.753[rgb]0.0, 0.0, 0.0 gemma3-27b Bulgarian 99.2(p=0.000)99.2~(p=0.000)
Gheg Albanian 80.4(p=0.050)80.4~(p=0.050)
Polish 98.8(p=0.002)98.8~(p=0.002)